What is the probability of rolling a 3 on two consecutive rolls of a fair die? Round your answer to the nearest hundredth. a.0.01 b 0.03 C.9.13 d.0.17
Answer by fractalier(6550)

Since the probability of rolling a 3 is 1/6, the probability of doing it twice in a row is 1/6 * 1/6 = 1/36 = 2.78% or .03
Choice B.